OpenStudy (laronjaxon):

solve for x 2/5+3/5x=x+5/10

OpenStudy (anonymous):

Put all x on the left side of the equation and all numbers on the right side. You have an equation like this\[\frac{ 2 }{ 5 }+\frac{3}{5}x=x+\frac{5}{10}\]Multiply the entire equation by 5 and you will get,\[2+3x=5x+\frac{5}{2}\]Now, put the x on the left side and the number on the right side, and you get,\[3x-5x=\frac{5}{2}-2\]This simplifies to,\[-2x=\frac{1}{2}\]Do the rest and get your solution.